OpenStudy (anonymous):

how would i find x from if the function f(x)=2 from he equation f(x)=1/x-1

OpenStudy (anonymous):

1/x-1=2 and solve for x or if you meant 1/(x-1)=2 don't know which one you meant

OpenStudy (anonymous):

i meant 1/(x-1)=2 .. i forgot how to.. idk.. lol i got as far as -1/(x-1) -2?

OpenStudy (anonymous):

Multiply (x-1) on both sides

OpenStudy (anonymous):

idk... my brain is so off today sorry

OpenStudy (anonymous):

i bet it is \[\frac{1}{x-1}=2\] \[1=2(x-1)\] \[1=2x-2\] \[3=2x\] \[x=\frac{3}{2}\] but i could be wrong

OpenStudy (anonymous):

wait so with 2=1/(x-1) i multiply both sides by x-1?

OpenStudy (anonymous):

OOOOH i got it.. -.- . wow .. that was so simple. kay thanks !